Find the inverse of the one to one function.
F(x)=4/5x-1
Thank you.

Answer by stanbon(57967)

Find the inverse of the one to one function.
F(x)=4/5x-1
---------------
y = (4/5)x -1
Interchange x and y to get:
x = (4/5)y -1
solve for y to get the inverse:
5x = 4y -5
4y = 5x + 5
y = (5/4)x + 5
That is the inverse.
==========================
Cheers,
Stan H.